OpenStudy (anonymous):

What is the equation of the line passing through the points (–25, 50) and (25, 50) in slope-intercept form?

OpenStudy (anonymous):

50-50/25-(-25)=0/50=0 it is parallel to x-axis

OpenStudy (anonymous):

huh??

OpenStudy (anonymous):

\[m=\frac{ y_{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

yeah help me

OpenStudy (anonymous):

Well\[m=\frac{ 50-50 }{ 25+25 }\] \[m=\frac{ 0 }{ 50 }\] \[m=0\]

OpenStudy (anonymous):

Slope is 0

OpenStudy (anonymous):

thats not none of the choices

OpenStudy (anonymous):

lol ill put 50

OpenStudy (anonymous):

can you help me with more

OpenStudy (anonymous):

I wasn't finished but sure I'll help you with more.

OpenStudy (anonymous):

okay thank you :)

OpenStudy (anonymous):

What is the slope-intercept form of a line that passes through points (2, 11) and (4, 17)?

OpenStudy (anonymous):

\[m=\frac{ 17-11 }{ 4-2 }\] \[m=\frac{ 6 }{ 2 }\] \[m=3\] So your slope is 3 slope-int form is \[y=mx+b\] Plug in one point and the slope \[y-y_{1}=m(x-x _{1)}\] \[y-11=3(x-2)\] \[y-11=3x-6\] \[y=3x+5\]

OpenStudy (anonymous):

awww thanks i got more

OpenStudy (anonymous):

What is the y-intercept of a line that has a slope of and passes through point (8, 3)?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

-5 I want to say. Did the work in my head. I can help you more in just a little because I need to take my baby a bath.

OpenStudy (anonymous):

okay np :)

OpenStudy (anonymous):

Okay I'm done.